OpenStudy (anonymous):

For the following transformation T either give its standard matrix (i.e. the matrix relative to the standard bases of the domain of T and the codomain of T), if T is linear, or find a counter-example that demonstrates T is not linear. (see question for full T)... help, please? :S

OpenStudy (anonymous):

T\[{ \left(\begin{matrix}x_1 \\ x_2\\\ x_3\\\ x_4\end{matrix}\right) }\] = \[\left(\begin{matrix}x_1 - 5 x_2 - 6x_3 + x_4 \\ 2x_2 - x_3 - x_4x_3\end{matrix}\right)\] (sorry about the formatting)

OpenStudy (zarkon):

is \[T\left[2{ \left(\begin{matrix}1 \\ 1\\\ 1\\\ 1\end{matrix}\right) }\right]=2T\left[{ \left(\begin{matrix}1 \\ 1\\\ 1\\\ 1\end{matrix}\right) }\right]\]

OpenStudy (anonymous):

... Yes?

OpenStudy (zarkon):

no

OpenStudy (anonymous):

oh :(

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

wait - ill give that an actual go, rather than just guessing...

OpenStudy (zarkon):

ok ;)

OpenStudy (anonymous):

OK, so the first (T(2[x_1 ... ])) is [-18, -2], and the other is [-22,0] ... or something..? :/

OpenStudy (zarkon):

-22 in not correct

OpenStudy (zarkon):

is not

OpenStudy (zarkon):

it should be -18

OpenStudy (zarkon):

the -2 and 0 are correct...and that makes it not linear

OpenStudy (zarkon):

\[x_4\cdot x_3\] is a problem

OpenStudy (anonymous):

In general for a transformation to fail to be linear it must fail to satisfy one of these two properties: \[T(u + v) = T(u) + T(v)\]\[T(av) = aT(v)\]

OpenStudy (anonymous):

As Zarkon demonstrated \(T(av) \neq aT(v)\)

OpenStudy (anonymous):

Cheers - don't know why I struggle so much with this stuff!! I'm a calculus major, and have done 2 years of engineering mathematics, but linear algebra just does my head in! Expect a few more questions as the semester goes on haha thanks!

OpenStudy (zarkon):

"calculus major" didn't know there was such a thing

OpenStudy (anonymous):

In Australia - might be different to wherever you are :)

OpenStudy (zarkon):

I'm in the USA